OpenStudy (anonymous):

In the simplest RC circuit consisting of two circuit elements, one capacitor (C) and one resistor (R) in series the equation of the circuit is stated to be C*dv(t)/dt + v(t)/R=0 why is the sign positive for the current over the resistance? Wouldn't that mean that the currents "collide" at any given point in the circuit?

OpenStudy (anonymous):

\[C\times \frac{dv(t)}{dt} + \frac{v(t)}{R}=0\]

OpenStudy (vincent-lyon.fr):

This is because your capacitor is discharging, which means that you implicitly write : \(i(t)=-\Large \frac{dq}{dt}\)